OpenStudy (anonymous):

If the sides of a square are lengthened by 6cm, the area become 144cm². Find the length of the side of the original square?

OpenStudy (anonymous):

6

OpenStudy (anonymous):

(x+6)^2 = 12^2 x + 6 = 12 x = 6